OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 2x - y = 8.

OpenStudy (anonymous):

x-intercept= 4 y-intercept= -8

OpenStudy (a1234):

To find the x-intercept, make y as 0. To find the y intercept, make x as 0. For the y-intercept: 2x - y = 8 2(0) - y = 8 -y = 8 or y = -8 For the x-intercept: 2x - y = 8 2x - 0 = 8 2x = 8 x = 4